OpenStudy (anonymous):

Solve the following equations by either factoring or completing the square: x^2-4x+3=0

OpenStudy (anonymous):

(x-3)(x-1)

OpenStudy (anonymous):

(x-3)(x-1)=0 x=3 or 1

OpenStudy (anonymous):

can you show me how you got that?

OpenStudy (anonymous):

do you know how to multiply the polynomials together?

OpenStudy (anonymous):

like (x+a)(x+b)

OpenStudy (anonymous):

yeah.

hero (hero):

Quick Solution: x^2-4x+3=0 x^2-3x-1x + 3 = 0 x(x-3)-1(x-3) = 0 (x-3)(x-1) = 0 x = 1,3 Long Solution: When given an equation of the form ax^2 + bx + c = 0 and asked to find x, perform the following general steps: 1. Find two numbers that multiply to get a*c, yet add to get b. In this case, find two numbers that multiply to get 3, yet add to get -4. The two numbers are -3 and -1. 2. Split the middle term, -4x, into two terms using the two numbers that you found: x^2 -3x -1x + 3 = 0 Notice that -3x - 1x = -4x 3. Factor the first two terms, then factor the last two terms such that the remaining terms are two identical binomials: x(x-3) -1(x-3) = 0 4. Factor out the remaining binomial to create the product of two binomials: (x-3)(x-1) = 0 5. Use the zero product property (ab = 0) to create two separate equations, then solve for x: x - 3 = 0 x - 1 = 0 x = 3 x = 1